OpenStudy (anonymous):

3-|8-2x| is greater than or equal to -1

OpenStudy (anonymous):

@AccessDenied help?

OpenStudy (anonymous):

you will have to get the absolute value by itself first

OpenStudy (anonymous):

then you have to decide if its even possible.

OpenStudy (anonymous):

it is possible.

OpenStudy (anonymous):

ok then. once the abs value is by itself make 2 inequalities

OpenStudy (anonymous):

on one use the current sign on the right. on the other use the opp sign and change the inequality around. does that make sense? is the equation editor working today?

OpenStudy (anonymous):

dont think so.. 3-8-2x

OpenStudy (anonymous):

-(3-8-2x)?

OpenStudy (anonymous):

you would do 3-{8-2x} >= -1 -{8-2x}>=-4 {8-2x}<=4 <this could be solution or {8-2x}>=-4 <this

OpenStudy (anonymous):

my brackets mean absolute value btw

OpenStudy (anonymous):

no thats wrong

OpenStudy (anonymous):

how is it wrong.

OpenStudy (anonymous):

the answer is [2,6]

OpenStudy (anonymous):

that is what i got.

OpenStudy (anonymous):

But, I don't know how.

OpenStudy (anonymous):

solve both of the linear inequalities.

OpenStudy (anonymous):

let me review both of them again. maybe i missed something

OpenStudy (anonymous):

8-2x<=4 or 8-2x>=-4 solve both of these.

OpenStudy (anonymous):

Yes, I got it now :) Thank you for helping me

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

I have two other that I'm stuck on that I've tried over and over and still cannot seem to get it right, can you help me with those?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1st. 4-2{2-5x}<=2 I got 2/5 and 6/5 but thats wrong. 2nd. 5{3x+1}+4<=4 I got 2 and 3, but thats wrong too.

OpenStudy (anonymous):

on the top one. tell me what your equation is once you have the absolute value by itself.

OpenStudy (anonymous):

2*{2-5x}<=2? 2*{2-5x}<=-2?

OpenStudy (anonymous):

no you need to get rid of that 2 on the left. and you just forgot about your minus sign there. it should be 4-2{2-5x}<=2 -2{2-5x}<=-2 {2-5x}>=1 2-5x>=1 or 2-5x<=-1

OpenStudy (anonymous):

see what is happening in each step? or do we need to break a step or two down?

OpenStudy (anonymous):

divided by -1 on each side to get rid of it right?

OpenStudy (anonymous):

sorry -2

OpenStudy (anonymous):

by -2 is right. after you subtract the 4

OpenStudy (anonymous):

and when you divide by negative you flip the inequality

OpenStudy (anonymous):

yes, okay then solving I got 1/5 and 3/5?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Yay! :) Okay, next one.

OpenStudy (anonymous):

5{3x+1}+4<=4

OpenStudy (anonymous):

divide by 5 on each side?

OpenStudy (anonymous):

this one is kind of interesting. what did you get once you got the abs by itself?

OpenStudy (anonymous):

{3x+1}+4<=4/5?

OpenStudy (anonymous):

3x+5<=4/5?

OpenStudy (anonymous):

you cant add something from outside the abs value. i would subtract the 4 first. then divide by the 5

OpenStudy (anonymous):

5{3x+1}=0?

OpenStudy (anonymous):

then divide both sides by 5

OpenStudy (anonymous):

and your putting an equal sign in there your still on inequalities right?

OpenStudy (anonymous):

{3x+1}<=0/5 and yes

OpenStudy (anonymous):

my bad sorry

OpenStudy (anonymous):

its good. so what do you get.

OpenStudy (anonymous):

-1/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome! Thank you :) Makes more sense now.

OpenStudy (anonymous):

no problem.